# A Spatial Gap in the Sky Distribution of Fast Radio Burst Detections Coinciding with Galactic Plasma Overdensities

### v1

@@ 3.1 Statistical Significance of the Gap
3.2 Effect of the Sky Temperature
@@
The spatial gap is towards Cygnus X (J. H. Piddington & H. C. Minnett, 1952b), a bright source of radio continuum at CHIME’s observing frequencies which could reduce sensitivity to a background burst. Using the Haslam 408 MHz radio sky temperature map (C. G. T. Haslam et al., 1982), scaled to the central frequency of the CHIME band with a scaling index of 2.6-2.6, the average sky temperature within the detection gap is $35$ K. To capture the effect of higher temperatures present in smaller sub-regions within the area, we use the 99th percentile $T_{\rm sky}=108$ K.
@@
follows directly from the standard radiometer equation as used in FRB and pulsar literature (e.g., J. M. Cordes & M. A. McLaughlin, 2003; C. Patel et al., 2018), where $S_{i}$ is the intrinsic flux density, $n_{p}$ is the number of polarizations measured, $\Delta\nu$ is the bandwidth and $\beta$ is the correction factor for digitization. Burst detectability for CHIME/FRB is governed by the intrinsic burst width $W_{i}$, intrinsic flux, DM and $\tau_{\rm sc}$. The observed peak flux density for an FRB decreases as its broadened pulse width - $W_{b}$ - increases, defined as (D. W. Gardenier et al., 2021)
@@
where $z_{0}$ is the redshift and $t_{\rm chan}$ depends linearly on the DM as described by Equation 8 in §4.2.
To compute the expected number of bursts within the detection gap, we perform simulations using subsamples of bursts from Catalog 2, explicitly accounting for the effect of elevated sky temperature. To ensure approximately uniform exposure and sensitivity, we select all Catalog 2 bursts in the similar declination range of $30\degree<\delta<50\degree$, resulting in 921 bursts. This sample serves as an empirical proxy for the intrinsic burst sample seen at this declination range in the absence of the influence of Cygnus X. From §3.1, the expected number of bursts originating behind the gap (i.e., before accounting for the elevated $T_{\rm sky}$ in the Cygnus X region) is 19, following a Poisson distribution. We generate 106 simulated realizations as follows:
@@
  • Randomly select $N_{\rm gap}$ bursts from the Catalog 2 subsample of 921 bursts, using their intrinsic burst width ($W_{i}$), sky temperature corresponding to their positions $T_{\rm sky,i}$ and signal-to-noise $(S/N)_{i}$ as burst parameters;
@@
4 Assessing Possible Origins for the Detection Gap
Figure 2: A comparison of literature values of scattering in and surrounding the Cygnus X region, as described in §4.1, overlaid on the Planck emission measure map (Planck Collaboration et al. 2016). The angular broadening measurements for quasars and Cygnus X-3 have been converted to scattering timescales using Equation 7, and scaled to 600 MHz using $\theta\propto\nu^{-2}$. Scattering measurements for pulsars and FRB 20210705 are extrapolated to 600 MHz using $\tau\propto\nu^{-4}$. This shows that sightlines through the Cygnus X region can produce $\sim 10-1,000\,$ms of temporal scattering in CHIME’s observing band, which would the hinder detection of an FRB (see §4.4). A white polygon encloses the area with zero FRB detections, formed by Delaunay triangulation as described in §3.
4.1 Background on the Cygnus X region
While Cygnus X is comprised of many distinct star forming regions, studies of molecular line observations have argued that they are associated, and at a similar distance $d_{\rm Cyg}$ (N. Schneider et al., 2006). This association was later supported using parallax measurements with masers (K. L. J. Rygl et al., 2012), and with Gaia (S. R. Berlanas et al., 2019), finding distance measurements spanning a range $1.3<d_{\rm Cyg}<1.7$ kpc. We adopt a value of $d_{\rm Cyg}=1.5$ kpc throughout our work (following discussion of F. Comerón et al. 2020), but emphasize that it does not strongly affect our analysis. At $d_{\rm Cyg}=1.5$ kpc, the $\sim 18\degree$ angular diameter of the spatial gap corresponds to $\sim 250\,$ pc. Additionally, the electron temperature $T_{\rm e}$ is a function of Galactocentric radius $R_{\rm gal}$, fit by C. Quireza et al. (2006) as $T_{\rm e}(\text{K})=(5780\pm 350)+(287\pm 46)R_{\rm Gal}.$ We adopt $T_{\rm e}=8000\,$K throughout for the Cygnus X region, appropriate for $R_{\rm Gal}\approx 8\,$kpc.
4.1.1 Emission Measure
From the Planck emission measure (EM) map (Planck Collaboration et al. 2016), which traces the free-free emission from ionized gas and has a resolution of $13.7^{\prime}\approx 3\,$pc, the maximum EM of the Cygnus X region is $9400$ $\text{ pc cm}^{-6}$. On smaller angular scales, the Cygnus X region contains many thin filamentary structures, which are overdense and overpressured compared to the average properties of their surroundings (K. L. Emig et al., 2022). The maximum EM in these filaments approaches $10^{5}$ $\text{ pc cm}^{-6}$ with a median value of 5200 $\text{ pc cm}^{-6}$, and electron densities of $n_{e}\approx 10-400$ $\text{ cm}^{-3}$, compared to the median of $n_{e}\approx 35$ $\text{ cm}^{-3}$. The density of the volume-filling ionized gas in the region is $n_{e}\approx 35$ $\text{ cm}^{-3}$, which itself is far above the volume-averaged plasma density in the WIM of $n_{e}\approx 0.03$ $\text{ cm}^{-3}$ (P. O. Mezger, 1978; K. L. Emig et al., 2022).
The observables $\text{DM}=\int_{0}^{L}n_{e}(l)dl$ and $\text{EM}=\int_{0}^{L}n_{e}(l)^{2}dl$ are physically related quantities where $n_{e}(l)$ is the free electron density (in  $\text{ cm}^{-3}$) at a distance $l$ (in pc) along the line of sight, and L is the total path length through the ionized medium (in pc). Following S. K. Ocker et al. (2024), we assume that a path length L with volume-average electron density $n_{e}$ is filled with dense cloudlets with a volume filling factor $f\leq 1$, and internal density $n_{e,i}$, such that $n_{e}=fn_{e,i}$. The other parameters are the variance of electron density within a cloudlet $\epsilon^{2}=\langle(\delta n_{e,i})^{2}\rangle/n_{e,i}^{2}\leq 1$, and between cloudlets $\zeta=\langle n_{e,i}^{2}\rangle/\langle n_{e,i}\rangle^{2}\geq 1$. Following this prescription, the DM and EM are related as
Based on the above, a source seen through Cygnus X at the maximum $\text{EM}\approx 10^{4}$ $\text{ pc cm}^{-6}$ (from the Planck map) with path length $\rm L\approx 500\,$pc could accrue $\text{DM}\sim 2200$ $\text{ pc cm}^{-3}$. As will be described in §4.1.3, §4.1.4, this estimate is in line with the highest DMs of pulsars in the region, and that of the background FRB 20210705 discovered at L-band with the Five-hundred-meter Aperture Spherical radio Telescope (FAST). As discussed in §4.1.1, the peak EM in filaments reaches $10^{5}$ $\text{ pc cm}^{-6}$, although most filaments have thickness on the order of a few pc (K. L. Emig et al., 2022). A sightline could, in principle, accumulate a larger DM if it is seen directly through the long axis of a filament, but this is unlikely, and not representative of an average line-of-sight through the Cygnus X region.
4.1.2 Angular Broadening
Background quasars, if intrinsically compact enough, can show signs of angular broadening when observed through multi-frequency very long baseline interferometry (VLBI), characterized by broadening $\theta\propto\nu^{\alpha}$, scaling with $\alpha\approx-2$. Early VLBI observations of background quasars show angular broadening with full-width at half-maximum (FWHM) of $\theta_{\rm FWHM}\approx 10-200\,$ mas referenced to 1 GHz within Galactic longitudes $60\degree\lesssim l\lesssim 80\degree$, greater than the surroundings (A. L. Fey et al., 1989, 1991). Observations of Cygnus X-3 (not associated with, but background to the Cygnus X region), show scatter broadening of 220250220-250\, mas at 1 GHz (L. A. Molnar et al., 1995). Many additional VLBI studies of quasars through Cygnus X (e.g. T. J. W. Lazio & A. L. Fey 2001; K. M. Desai & A. L. Fey 2001; K. É. Gabányi et al. 2006) show angular broadening of a similar order, as well as extreme scattering event-like structures seen in the light-curve of quasar B2005+403, implying $\sim 0.7\,$AU plasma structures (T. A. Koryukova et al., 2023).
Angular broadening, combined with the known distance to Cygnus X, can be used to predict the scattering time delay. For a scatter-broadened image with $\theta_{\rm FWHM}$, the scattering timescale $\tau_{\rm sc}$, defined as the $1/e$ timescale for a decaying exponential, is related to $\theta_{\rm FWHM}$ as (e.g. J. M. Cordes & S. Chatterjee 2019):
$\tau_{\rm sc}\approx\frac{d_{\rm eff}\theta_{\rm FWHM}^{2}}{8\ln(2)c},$ (7)
4.1.3 Pulsars
There are >90>90 known pulsars within our detection gap (R. N. Manchester et al., 2005). With surveys from more sensitive telescopes, many newly discovered pulsars in this region have DMs greater than the full expected Milky Way contribution from Galactic-$n_{e}$ models (J. L. Han et al., 2025), as the abundance of plasma within Cygnus X is not properly accounted for. Without a reliable $n_{e}$ model along these sightlines, and with few independent distance measurements, it is difficult to know which pulsars are foreground or background to the star-forming complex. As the complex itself is likely the cause of the discrepancy, the sources with DMs far exceeding model predictions can be assumed to be within or behind Cygnus X.
The FAST telescope recently measured scattering of 17 pulsars within our detection gap, spanning a range of $\tau_{\rm sc,\,1\textrm{ GHz}}\approx 1-300$ ms, and DMs 200950200-950 $\text{ pc cm}^{-3}$ (W. C. Jing et al., 2025). Other pulsars with scattering measurements include PSR J2108+5001 discovered by the CHIME All-sky Multi-day Pulsar Stacking Search (The CHAMPSS Collaboration et al. 2025) located just outside of our detection gap boundary, and 9 pulsars discovered with the Green Bank Telescope (GBT) at 820 MHz (A. E. McEwen et al., 2024).
4.1.4 FRBs at Higher Observing Frequencies
[…]
4.2 Biases from DM Models and Dispersion Smearing
As described in §2.3, the CHIME/FRB backend classifies signals as Galactic, ambiguous, or extragalactic based on their position and DM, with extragalactic FRBs identified when their DMs exceed the maximum Galactic contributions predicted by the NE2001 and YMW16 models. Structural complexity of the ISM in the Milky Way plane leads to greater uncertainty in dispersion and scattering predicted by these models at lower Galactic latitudes compared to lines of sight at higher Galactic latitudes. In principle, these uncertainties could lead to misclassifications of FRBs as instead arising from Galactic sources due to an overestimation of the Galactic DM along the LOS by available models. The FRB 20220319D, detected and precisely localized at a distance of 50 Mpc by the Deep Synoptic Array (DSA-110), is one such example. It exhibits a DM (110.95 $\text{ pc cm}^{-3}$) lower than those predicted by the NE2001 (132.9 $\text{ pc cm}^{-3}$) and YMW16 (187.7 $\text{ pc cm}^{-3}$) models (V. Ravi et al., 2025) for extragalactic objects along its LOS ($l=129.2\degree,\ b=9.1\degree$). This highlights how overestimated model DMs may hinder the detectability of nearby, low-DM FRBs in pipelines with DM-based filtering. As discussed in §4.1.3, studies of pulsars in the background of Cygnux X have shown that the Galactic-$n_{e}$ models instead make it more likely that unknown pulsars in the detection gap with unusually high DMs might be misclassified as FRBs.
We acknowledge the possibility that the Cygnus X complex is so poorly accounted for in the models that they could underestimate the maximum DM by an order of magnitude, i.e., that the true value of ${\rm DM}_{\rm max}\geq 5000\textrm{ pc cm}^{-3}$ along the LOS. At such high DMs, intra-channel dispersion smearing could partially emulate scatter-broadening and reduce pulse brightness at low frequencies. However, we believe that DM smearing can be ruled out as the origin of the detection gap. Temporal broadening due to intra-channel dispersion smearing (J. M. Cordes & M. A. McLaughlin, 2003) is given by:
The CHIME/FRB detection pipeline searches to a maximum DM of 13000 $\text{ pc cm}^{-3}$ (CHIME/FRB Collaboration et al. 2018). In order to produce $\rm t_{\rm chan}\approx 8$ ms at 600 MHz – comparable to the mean value of $\tau_{\rm sc}$ measured in Catalog 2 (1 ms at 1GHz; see §4.1.4) – a value of ${\rm DM}>8000\textrm{ pc cm}^{-3}$ is required. This estimate is nearly an order of magnitude higher than the DM inferred from the maximum EM in the Cygnus X region as discussed in §4.1.1.
4.3 Free-free Absorption in the Direction of Cygnus X
Here, we examine whether the FRB detection gap towards Cygnus X could be caused by free-free absorption. The free-free opacity is well approximated by (P. G. Mezger & A. P. Henderson, 1967)
As discussed in §4.1, the Planck EM map for the Cygnus X region indicates that $\rm{EM}\approx 9400$ $\text{ pc cm}^{-6}$, while detailed studies of overdense and overpressured filaments show peak values of $\rm{EM}\approx 10^{5}$ $\text{ pc cm}^{-6}$ K. L. Emig et al. (2022).
Inserting $\rm{EM}\approx 10^{4}$ $\text{ pc cm}^{-6}$ for the Cygnus X region (S. K. Ocker et al., 2024), and $T_{e}\approx 8000\,$K, the opacity is $\tau_{ff}\approx 0.02$ at $\nu=400\,$MHz, the bottom of our band. For the maximum measured emission measures of $\rm{EM}\approx 10^{5}$ $\text{ pc cm}^{-6}$, free-free absorption begins to matter, with $\tau_{ff}\approx 0.2$ at $\nu=400\,$MHz, but still a small effect, and with $\tau_{ff}<0.1$ averaged across the band. The detection of background quasars at $350$ MHz (e.g. T. J. W. Lazio & A. L. Fey 2001), and prevalence of free-free emission at 148148\,MHz (e.g. K. L. Emig et al. 2022) also qualitatively suggests that free-free absorption is not dominating at low radio frequencies in this region. We therefore conclude that while free-free absorption could have a small effect on a background FRB, it is not the primary cause of the detection gap.
4.4 Empirical Limits on Scattering Timescale

### v2

@@ III Statistical Significance of the Detection Gap
IV Assessing Possible Origins for the Detection Gap
Figure 3: A comparison of literature values of scattering in and surrounding the Cygnus X region, as described in §IV.1, overlaid on the Planck emission measure map (planckforegrounds16). The angular broadening measurements for quasars and Cygnus X-3 have been converted to scattering timescales using Equation 5, and scaled to 600 MHz using $\theta\propto\nu^{-2}$. Scattering measurements for pulsars and FRB 20210705 are extrapolated to 600 MHz using $\tau\propto\nu^{-4}$. This shows that sightlines through the Cygnus X region can produce $\sim 10-1,000\,$ms of temporal scattering in CHIME’s observing band, which would the hinder detection of an FRB (see §IV.5). A white polygon encloses the area with zero FRB detections, formed by Delaunay triangulation as described in §III.
IV.1 Background on the Cygnus X region
While Cygnus X is comprised of many distinct star forming regions, studies of molecular line observations have argued that they are associated, and at a similar distance $d_{\rm Cyg}$ (schneider+06). This association was later supported using parallax measurements with masers (rygl+12), and with Gaia (berlanas+19), finding distance measurements spanning a range $1.3<d_{\rm Cyg}<1.7$ kpc. We adopt a value of $d_{\rm Cyg}=1.5$ kpc throughout our work (following discussion of comeron+20), but emphasize that it does not strongly affect our analysis. At $d_{\rm Cyg}=1.5$ kpc, the $\sim 18\degree$ angular diameter of the spatial gap corresponds to $\sim 250\,$ pc. Additionally, the electron temperature $T_{\rm e}$ is a function of Galactocentric radius $R_{\rm gal}$, fit by quireza+06 as $T_{\rm e}(\text{K})=(5780\pm 350)+(287\pm 46)R_{\rm Gal}.$ We adopt $T_{\rm e}=8000\,$K throughout for the Cygnus X region, appropriate for $R_{\rm Gal}\approx 8\,$kpc.
IV.1.1 Emission Measure
From the Planck emission measure (EM) map (planckforegrounds16), which traces the free-free emission from ionized gas and has a resolution of $13.7^{\prime}\approx 3\,$pc, the maximum EM of the Cygnus X region is $9400$ $\text{ pc cm}^{-6}$. On smaller angular scales, the Cygnus X region contains many thin filamentary structures, which are overdense and overpressured compared to the average properties of their surroundings (emig+22). The maximum EM in these filaments approaches $10^{5}$ $\text{ pc cm}^{-6}$ with a median value of 5200 $\text{ pc cm}^{-6}$, and electron densities of $n_{e}\approx 10-400$ $\text{ cm}^{-3}$, compared to the median of $n_{e}\approx 35$ $\text{ cm}^{-3}$. The density of the volume-filling ionized gas in the region is $n_{e}\approx 35$ $\text{ cm}^{-3}$, which itself is far above the volume-averaged plasma density in the WIM of $n_{e}\approx 0.03$ $\text{ cm}^{-3}$ (mezger+78; emig+22).
The observables $\text{DM}=\int_{0}^{L}n_{e}(l)dl$ and $\text{EM}=\int_{0}^{L}n_{e}(l)^{2}dl$ are physically related quantities where $n_{e}(l)$ is the free electron density (in  $\text{ cm}^{-3}$) at a distance $l$ (in pc) along the line of sight, and L is the total path length through the ionized medium (in pc). Following OAL+24, we assume that a path length L with volume-average electron density $n_{e}$ is filled with dense cloudlets with a volume filling factor $f\leq 1$, and internal density $n_{e,i}$, such that $n_{e}=fn_{e,i}$. The other parameters are the variance of electron density within a cloudlet $\epsilon^{2}=\langle(\delta n_{e,i})^{2}\rangle/n_{e,i}^{2}\leq 1$, and between cloudlets $\zeta=\langle n_{e,i}^{2}\rangle/\langle n_{e,i}\rangle^{2}\geq 1$. Following this prescription, the DM and EM are related as
Based on the above, a source seen through Cygnus X at the maximum $\text{EM}\approx 10^{4}$ $\text{ pc cm}^{-6}$ (from the Planck map) with path length $\rm L\approx 500\,$pc could accrue $\text{DM}\lesssim 2200$ $\text{ pc cm}^{-3}$. As will be described in §IV.1.3, §IV.1.4, this estimate is in line with the highest DMs of pulsars in the region, and that of the background FRB 20210705 discovered at L-band with the Five-hundred-meter Aperture Spherical radio Telescope (FAST). As discussed in §IV.1.1, the peak EM in filaments reaches $10^{5}$ $\text{ pc cm}^{-6}$, although most filaments have thickness on the order of a few pc (emig+22). A sightline could, in principle, accumulate a larger DM if it is seen directly through the long axis of a filament, but this is unlikely, and not representative of an average line-of-sight through the Cygnus X region.
IV.1.2 Angular Broadening
Background quasars, if intrinsically compact enough, can show signs of angular broadening when observed through multi-frequency very long baseline interferometry (VLBI), characterized by broadening $\theta\propto\nu^{\alpha}$, scaling with $\alpha\approx-2$. Early VLBI observations of background quasars show angular broadening with full-width at half-maximum (FWHM) of $\theta_{\rm FWHM}\approx 10-200\,$ mas referenced to 1 GHz within Galactic longitudes $60\degree\lesssim l\lesssim 80\degree$, greater than the surroundings (fey+89; fey+91). Observations of Cygnus X-3 (not associated with, but background to the Cygnus X region), show scatter broadening of 220250220-250\, mas at 1 GHz (CygnusX-molnar). Many additional VLBI studies of quasars through Cygnus X (e.g. lazio+01; desai+01; gabanyi+06) show angular broadening of a similar order, as well as extreme scattering event-like structures seen in the light-curve of quasar B2005+403, implying $\sim 0.7\,$AU plasma structures (koryukova+23).
Angular broadening, combined with the known distance to Cygnus X, can be used to predict the scattering time delay. For a scatter-broadened image with $\theta_{\rm FWHM}$, the scattering timescale $\tau_{\rm sc}$, defined as the $1/e$ timescale for a decaying exponential, is related to $\theta_{\rm FWHM}$ as (e.g. cc19):
$\tau_{\rm sc}\approx\frac{d_{\rm eff}\theta_{\rm FWHM}^{2}}{8\ln(2)c},$ (5)
IV.1.3 Pulsars
There are >90>90 known pulsars within our detection gap (manchester+05). With surveys from more sensitive telescopes, many newly discovered pulsars in this region have DMs greater than the full expected Milky Way contribution from Galactic-$n_{e}$ models (han+25), as the abundance of plasma within Cygnus X is not properly accounted for. Without a reliable $n_{e}$ model along these sightlines, and with few independent distance measurements, it is difficult to know which pulsars are foreground or background to the star-forming complex. As the complex itself is likely the cause of the discrepancy, the sources with DMs far exceeding model predictions can be assumed to be within or behind Cygnus X.
The FAST telescope recently measured scattering of 17 pulsars within our detection gap, spanning a range of $\tau_{\rm sc,\,1\textrm{ GHz}}\approx 1-300$ ms, and DMs 200950200-950 $\text{ pc cm}^{-3}$ (jing+25). Other pulsars with scattering measurements include PSR J2108+5001 discovered by the CHIME All-sky Multi-day Pulsar Stacking Search (CHAMPSS+25) located just outside of our detection gap boundary, and 9 pulsars discovered with the Green Bank Telescope (GBT) at 820 MHz (mcewan+24).
IV.1.4 FRBs at Higher Observing Frequencies
[…]
IV.2 Biases from DM Models and Dispersion Smearing
As described in §II.3, the CHIME/FRB pipeline uses Galactic DM predictions from NE2001 and YMW16 to classify events as Galactic or extragalactic. This raises a general concern that DM-based thresholds could cause nearby, low-DM FRBs to be missed, if the models overestimate the Galactic DM along the LOS. For example, FRB 20220319D (DM = 110.95 $\text{ pc cm}^{-3}$, distance = 50 Mpc) detected by the Deep Synoptic Array (DSA-110), exhibits a DM lower than the model predictions (rcc+25). However, in the Cygnus X region, the situation is reversed and the Galactic-$n_{e}$ models underestimate the DM (see pulsars in §IV.1.3). As a result, pulsars with unusually high DMs are more likely to be misclassified as FRBs, not vice versa. Therefore, this classification bias cannot explain the absence of FRBs in the detection gap.
We acknowledge the possibility that the Cygnus X complex is so poorly accounted for in the models that they could underestimate the maximum DM by an order of magnitude, i.e., that the true value of ${\rm DM}_{\rm max}\geq 5000\textrm{ pc cm}^{-3}$ along the LOS. At such high DMs, intra-channel dispersion smearing could partially emulate scatter-broadening and reduce pulse brightness at low frequencies. However, we believe that DM smearing can be ruled out as the origin of the detection gap. Temporal broadening due to intra-channel dispersion smearing (cm03) is given by:
The CHIME/FRB detection pipeline searches to a maximum DM of 13000 $\text{ pc cm}^{-3}$ (abb+18). In order to produce $\rm t_{\rm chan}\approx 8$ ms at 600 MHz – comparable to the mean value of $\tau_{\rm sc}$ measured in Catalog 2 (1 ms at 1GHz; see §IV.1.4) – a value of ${\rm DM}>8000\textrm{ pc cm}^{-3}$ is required. This estimate is substantially higher than the DM inferred from the maximum EM in the Cygnus X region as discussed in §IV.1.1.
IV.3 Free-free Absorption in the Direction of Cygnus X
Here, we examine whether the FRB detection gap towards Cygnus X could be caused by free-free absorption. The free-free opacity is well approximated by (mezger67)
As discussed in §IV.1, the Planck EM map for the Cygnus X region indicates that $\rm{EM}\approx 9400$ $\text{ pc cm}^{-6}$, while detailed studies of overdense and overpressured filaments show peak values of $\rm{EM}\approx 10^{5}$ $\text{ pc cm}^{-6}$ emig+22.
Inserting $\rm{EM}\approx 10^{4}$ $\text{ pc cm}^{-6}$ for the Cygnus X region (OAL+24), and $T_{e}\approx 8000\,$K, the opacity is $\tau_{ff}\approx 0.02$ at $\nu=400\,$MHz, the bottom of our band. For the maximum measured emission measures of $\rm{EM}\approx 10^{5}$ $\text{ pc cm}^{-6}$, free-free absorption begins to matter, with $\tau_{ff}\approx 0.2$ at $\nu=400\,$MHz, but still a small effect, and with $\tau_{ff}<0.1$ averaged across the band. The detection of background quasars at $350$ MHz (e.g. lazio+01), and prevalence of free-free emission at 148148\,MHz (e.g. emig+22) also qualitatively suggests that free-free absorption is not dominating at low radio frequencies in this region. We therefore conclude that while free-free absorption could have a small effect on a background FRB, it is not the primary cause of the detection gap.
IV.4 Effect of the Sky Temperature
@@
The spatial gap is towards Cygnus X (piddington+52), a bright source of radio continuum at CHIME’s observing frequencies which could reduce sensitivity to a background burst. Using the Haslam 408 MHz radio sky temperature map (haslam), scaled to the central frequency of the CHIME band with a scaling index of 2.6-2.6, the average sky temperature within the detection gap is $35$ K. To capture the effect of higher temperatures present in smaller sub-regions within the area, we use the 99th percentile $T_{\rm sky}=108$ K.
@@
follows directly from the standard radiometer equation as used in FRB and pulsar literature (e.g., cm03; pab+18), where $S_{i}$ is the intrinsic flux density, $n_{p}$ is the number of polarizations measured, $\Delta\nu$ is the bandwidth and $\beta$ is the correction factor for digitization. Burst detectability for CHIME/FRB is governed by the intrinsic burst width $W_{i}$, intrinsic flux, DM and $\tau_{\rm sc}$. The observed peak flux density for an FRB decreases as its broadened pulse width - $W_{b}$ - increases, defined as (gcl+21)
@@
where $z_{0}$ is the redshift and $t_{\rm chan}$ depends linearly on the DM as described by Equation 6 in §IV.2.
To compute the expected number of bursts within the detection gap, we perform simulations using subsamples of bursts from Catalog 2, explicitly accounting for the effect of elevated sky temperature. To ensure approximately uniform exposure and sensitivity, we select all Catalog 2 bursts in the similar declination range of $30\degree<\delta<50\degree$, resulting in 919 bursts. This sample serves as an empirical proxy for the intrinsic burst sample seen at this declination range in the absence of the influence of Cygnus X. From §III, the expected number of bursts originating behind the gap (i.e., before accounting for the elevated $T_{\rm sky}$ in the Cygnus X region) is 19, following a Poisson distribution. We generate 106 simulated realizations as follows:
@@
  • Randomly select $N_{\rm gap}$ bursts from the Catalog 2 subsample of 919 bursts, using their intrinsic burst width ($W_{i}$), sky temperature corresponding to their positions $T_{\rm sky,i}$ and signal-to-noise $(S/N)_{i}$ as burst parameters;
@@
IV.5 Empirical Limits on Scattering Timescale